# Universal quantum control over Majorana zero modes

### v1

@@ Acknowledgments
We acknowledge grant support from the Science and Technology Program of Zhejiang Province (No. 2025C01028).
@@ Appendix A UQC theory in parametric space
This appendix extends the universal quantum control theory Jin and Jing (2025a, b, c, d) from the time domain to the parametric space. Consider a $K$-dimensional system under a general effective Hamiltonian $H_{\rm eff}(\vec{R})$ with a varying control parameter vector $\vec{R}\equiv(R_{1},R_{2},\cdots,R_{M})$. $\dot{\vec{R}}\equiv d\vec{R}/dt=0$ means adiabatic condition. $H_{\rm eff}(\vec{R})$ can be obtained by such as Eq. (8), exhibiting an explicit or implicit time dependence, i.e., $R_{j}\equiv R_{j}(t)$ with $1\leq j\leq M$. The system dynamics is described by the parameter-dependent Schrödinger equation
$i\frac{d|\psi_{n}(\vec{R})\rangle}{dt}=H_{\rm eff}(\vec{R})|\psi_{n}(\vec{R})\rangle,$ (35)
@@
In UQC theory Jin and Jing (2025a, b, c), it is convenient to treat the system dynamics within an ancillary picture spanned by a completed and orthonormal set of ancillary basis states, i.e., $|\mu_{k}(\vec{R})\rangle$’s, $1\leq k\leq K$. $|\mu_{k}(\vec{R})\rangle$’s span the same Hilbert space as $|\psi_{n}(\vec{R})\rangle$’s. In the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, where $\vec{R}_{0}$ is a fixed parameter vector representing but not limited to the boundary condition $\vec{R}(t=t_{0})$, the Hamiltonian can be transformed as
$H_{\rm rot}(\vec{R})=V^{\dagger}(\vec{R})H_{\rm eff}(\vec{R})V(\vec{R})-iV^{\dagger}(\vec{R})\dot{V}(\vec{R})$ (36)
@@
$-i\dot{\vec{R}}\cdot\langle\mu_{k}(\vec{R})|\nabla_{\vec{R}}|\mu_{n}(\vec{R})\rangle\Big{]}|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|$
@@
where the dynamical component $\mathcal{H}_{kn}(\vec{R})$ and the gauge-potential component $\mathcal{A}_{kn}(\vec{R})$ Kolodrubetz et al. (2017); Claeys et al. (2019); Takahashi and del Campo (2024) constitute the elements of the matrixes $\mathcal{H}(\vec{R})$ and $\mathcal{A}(\vec{R})$, respectively, at the $k$th row and the $n$th column. The purely geometrical matrix $\mathcal{A}(\vec{R})$ relates uniquely to the manifold geometry determined by the ancillary basis states $|\mu_{k}(\vec{R})\rangle$’s. The Schrödinger equation (35) can then be transformed as
$i\frac{d|\psi_{n}(\vec{R})\rangle_{\rm rot}}{dt}=H_{\rm rot}(\vec{R})|\psi_{n}(\vec{R})\rangle_{\rm rot}$ (37)
@@
The parameter-dependent Schrödinger equation (37) can be exactly solved when $H_{\rm rot}(\vec{R})$ is diagonalizable Jin and Jing (2025a, b, c), particularly, when Eq. (36) can be simplified as
@@
Transformed back to the original picture via Eq. (38), the evolution operator in the original picture can be written as
@@
Due to the theory of UQC, the diagonalization condition of $H_{\rm rot}(\vec{R})$ (36) in the $k$th ancillary basis state $|\mu_{k}(\vec{R})\rangle$ is given by
@@
with the projection operator $\Pi_{k}(\vec{R})\equiv|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|$. It admits a compact form as
@@
$=$ $i\dot{\vec{R}}\cdot\Big{[}\nabla_{\vec{R}}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|+|\mu_{k}(\vec{R})\rangle\langle\mu_{k}{(\vec{R})}|\nabla_{\vec{R}}\Big{]}$
@@
where the second equivalence used the relation for the gauge field, i.e., $\mathcal{A}(\vec{R})|\mu_{k}(\vec{R})\rangle=i\sum_{j=1}^{M}\dot{R}_{j}|\partial_{R_{j}}\mu_{k}(\vec{R})\rangle$, as defined in Eq. (36). Equation (44) can be recast as a von Neumann-type equation Jin and Jing (2025a, b, c):
$\frac{d}{dt}\Pi_{k}(\vec{R})=-i\left[H_{\rm eff}(\vec{R}),\Pi_{k}(\vec{R})\right].$ (45)
When $k$ in Eq. (45) runs from 1 to $K$, $H_{\rm rot}(\vec{R})$ can be completely diagonalized with a full rank.
@@ Appendix B UQC theory in parametric space with error correction
This appendix provides a recipe for the universal quantum control theory in parametric space to suppress the systematic errors Jin and Jing (2025c). When unavoidable fluctuations appear in the control parameters, the system dynamics is actually subject to the full Hamiltonian $H(\vec{R})=H_{\rm eff}(\vec{R})+\epsilon H_{1}(\vec{R})$. Therefore, the parameter-dependent Schrödinger equation (35) becomes
$i\frac{d|\psi_{n}(\vec{R})\rangle}{dt}=H(\vec{R})|\psi_{n}(\vec{R})\rangle,$ (46)
where $H_{1}(\vec{R})$ is the error Hamiltonian and $\epsilon$ is a perturbative coefficient.
Consider the system dynamics in the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, the ideal rotated Hamiltonian in Eq. (39) can be perturbed by the error Hamiltonian as
@@
To identify and cancel the adverse effects induced by systematic errors, the system dynamics can be considered in the double-rotated picture with respect to $U_{\rm rot}(\vec{R})$ in Eq. (40). In particular, we have
$H_{\rm err}(\vec{R})$ $=U_{\rm rot}^{\dagger}(\vec{R})H_{\rm rot}(\vec{R})U_{\rm rot}(\vec{R})-iU_{\rm rot}^{\dagger}(\vec{R})\dot{U}_{\rm rot}(\vec{R})$ (48)
@@
Using the Magnus expansion Blanes et al. (2009), the evolution operator for $H_{\rm err}(\vec{R})$ in Eq. (48) can be expanded as (see Ref. Jin and Jing (2025c) for details)
@@
To evaluate the impact of the systematic errors, one can define the overlap or fidelity between the instantaneous state governed by $U_{\rm err}(\vec{R})$ in Eq. (50) and the target passage $|\mu_{k}(\vec{R}_{0})\rangle$ in the rotating frame Jin and Jing (2025c) as
$\mathcal{F}$ $\equiv|\langle\mu_{k}(\vec{R}_{0})|U_{\rm err}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle|^{2}$ (51)
@@
The error up to the second-order can be cancelled under the error correction mechanism Jin and Jing (2025c)
@@
This result is supported by the nonperturbative dynamical decoupling method Jing et al. (2013, 2015) and the Riemann-Lebesgue Lemma in mathematics.
@@
$\left|\frac{\partial}{\partial R}[f_{k}(R)-f_{n}(R)]\right|\gg\left|\frac{\partial\langle\mu_{k}(R)|H_{1}(R)|\mu_{n}(R)\rangle}{\partial R}\right|.$ (53)
Also, the gradient inequality with respect to the control parameters in Eq. (52) can be reformulated as a gradient inequality with respect to time:
$\left|\frac{\partial}{\partial t}[f_{k}(t)-f_{n}(t)]\right|\gg\frac{\partial}{\partial t}\left[\langle\mu_{k}(t)|H_{1}(t)|\mu_{n}(t)\rangle\right],$ (54)
which is consistent with that in Ref. Jin and Jing (2025c).

### v2

@@ Acknowledgments
We acknowledge grant support from the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
@@ Appendix A UQC theory in parametric space
This appendix extends the universal quantum control theory Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr; Jin2025Rydberg from the time domain to the parametric space. Consider a $K$-dimensional system under a general effective Hamiltonian $H_{\rm eff}(\vec{R})$ with a varying control parameter-vector $\vec{R}\equiv(R_{1},R_{2},\cdots,R_{L})$, exhibiting an explicit or implicit time dependence, i.e., $R_{j}\equiv R_{j}(t)$ with $1\leq j\leq L$. $\dot{\vec{R}}\equiv d\vec{R}/dt\approx 0$ means the adiabatic limit, by which the time evolution of the system is driven out of equilibrium at an infinitely slow rate. Our theory is however nonadiabatic. $H_{\rm eff}(\vec{R})$ can be obtained by such as Eq. (8). The system dynamics is described by the parameter-dependent Schrödinger equation
$i\frac{\partial|\psi_{n}(\vec{R})\rangle}{\partial t}=H_{\rm eff}(\vec{R})|\psi_{n}(\vec{R})\rangle,$ (40)
@@
In UQC theory Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr, it is convenient to treat the system dynamics within an ancillary picture spanned by a completed and orthonormal set of ancillary basis states, i.e., $|\mu_{k}(\vec{R})\rangle$’s, $1\leq k\leq K$. $|\mu_{k}(\vec{R})\rangle$’s span the same Hilbert space as $|\psi_{n}(\vec{R})\rangle$’s. In the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, where $\vec{R}_{0}$ is a fixed parameter vector representing but not limited to some boundary condition $\vec{R}(t=t_{0})$, the Hamiltonian can be transformed as
$H_{\rm rot}(\vec{R})=V^{\dagger}(\vec{R})H_{\rm eff}(\vec{R})V(\vec{R})-iV^{\dagger}(\vec{R})\partial_{t}V(\vec{R})$ (41)
@@
where the dynamical component $\mathcal{H}_{kn}(\vec{R})$ and the gauge-potential component $\mathcal{A}_{kn}(\vec{R})$ Michael2017Geometry; Claeys2019Floquet; Takahashi2024Shortcuts constitute the elements of the matrices $\mathcal{H}(\vec{R})$ and $\mathcal{A}(\vec{R})$, respectively, at the $k$th row and the $n$th column. The purely geometrical matrix $\mathcal{A}(\vec{R})$ relates uniquely to the manifold geometry determined by the ancillary basis states $|\mu_{k}(\vec{R})\rangle$’s, whose Berry connection is $i\langle\mu_{k}(\vec{R})|\nabla_{\vec{R}}|\mu_{n}(\vec{R})\rangle$. The Schrödinger equation (40) can then be transformed as
$i\frac{\partial|\psi_{n}(\vec{R})\rangle_{\rm rot}}{\partial t}=H_{\rm rot}(\vec{R})|\psi_{n}(\vec{R})\rangle_{\rm rot}$ (42)
@@
The parameter-dependent Schrödinger equation (42) can be exactly solved when $H_{\rm rot}(\vec{R})$ is diagonalizable Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr, or more precisely, when Eq. (41) can be simplified as
@@
Transformed back to the original picture via Eq. (43), the evolution operator can be written as
@@
Due to the theory of UQC, the diagonalization condition for $H_{\rm rot}(\vec{R})$ (41) in the $k$th ancillary basis state $|\mu_{k}(\vec{R})\rangle$ is given by
@@
with the projector $\Pi_{k}(\vec{R})\equiv|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|$. It admits a compact form as
@@
$=$ $i\dot{\vec{R}}\cdot\left[|\nabla_{\vec{R}}\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|+|\mu_{k}(\vec{R})\rangle\langle\nabla_{\vec{R}}\mu_{k}{(\vec{R})}|\right]$
@@
where the second equivalence used the relation for the gauge field, i.e., $\mathcal{A}(\vec{R})|\mu_{k}(\vec{R})\rangle=i\sum_{j=1}^{M}\dot{R}_{j}|\partial_{R_{j}}\mu_{k}(\vec{R})\rangle$, as defined in Eq. (41). Equation (49) can be recast as a von Neumann-type equation Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr:
$\frac{\partial}{\partial t}\Pi_{k}(\vec{R})=-i\left[H_{\rm eff}(\vec{R}),\Pi_{k}(\vec{R})\right].$ (50)
When $k$ in Eq. (50) runs from 1 to $K$, $H_{\rm rot}(\vec{R})$ can be completely diagonalized.
@@ Appendix B UQC theory in parametric space with error correction
This appendix provides a recipe for the universal quantum control theory in parametric space to suppress the systematic errors Jin2025ErrCorr. When the control parameters are subject to unavoidable fluctuations or unwanted crosstalk with the external degrees of freedom, the system dynamics is actually described by the realistic Hamiltonian $H(\vec{R})=H_{\rm eff}(\vec{R})+\epsilon H_{1}(\vec{R})$, where $H_{1}(\vec{R})$ is the error Hamiltonian and $\epsilon$ is a perturbative coefficient. Therefore, the parameter-dependent Schrödinger equation (40) becomes
$i\frac{\partial|\psi_{n}(\vec{R})\rangle}{\partial t}=H(\vec{R})|\psi_{n}(\vec{R})\rangle.$ (51)
Consider the system dynamics in the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, the ideal rotated Hamiltonian in Eq. (44) can be perturbed as
@@
To identify and neutralize the adverse effects induced by systematic errors, the system dynamics can be analyzed in the double-rotated picture with respect to $U_{\rm rot}(\vec{R})$ in Eq. (45). In particular, we have
$H_{\rm err}(\vec{R})$ $=U_{\rm rot}^{\dagger}(\vec{R})H_{\rm rot}(\vec{R})U_{\rm rot}(\vec{R})-iU_{\rm rot}^{\dagger}(\vec{R})\partial_{t}U_{\rm rot}(\vec{R})$ (53)
@@
Using the Magnus expansion Blanes2009Magnus, the evolution operator for $H_{\rm err}(\vec{R})$ in Eq. (53) can be expanded as
@@
To evaluate the impact of the systematic errors, one can define the overlap or fidelity between the instantaneous state governed by $U_{\rm err}(\vec{R})$ in Eq. (55) and the target passage $|\mu_{k}(\vec{R}_{0})\rangle$ in the rotating frame Jin2025ErrCorr as
$\mathcal{F}$ $\equiv\left|\langle\mu_{k}(\vec{R}_{0})|U_{\rm err}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle\right|^{2}$ (56)
@@
The error up to the second order of $\epsilon$ can be canceled under the error correction mechanism Jin2025ErrCorr
@@
This result is supported by the nonperturbative dynamical decoupling method Jing2013Nonperturbative; Jing2015Nonperturbative and the Riemann-Lebesgue lemma in mathematics.
@@
$\left|\frac{d}{dR}[f_{k}(R)-f_{n}(R)]\right|\gg\left|\frac{d\langle\mu_{k}(R)|H_{1}(R)|\mu_{n}(R)\rangle}{dR}\right|.$ (58)
Also, the gradient inequality with respect to the control parameter can be reformulated as a gradient inequality with respect to time:
$\left|\frac{d}{dt}[f_{k}(t)-f_{n}(t)]\right|\gg\frac{d}{dt}\left[\langle\mu_{k}(t)|H_{1}(t)|\mu_{n}(t)\rangle\right],$ (59)
which has been proposed in Ref. Jin2025ErrCorr.